OpenStudy (anonymous):

an algebraic expression that gives the area of a circle with a circumference ? Hint: The circumference of a circle with radius R is 2piR , and its area is piR^2

OpenStudy (anonymous):

Are you looking for the area of an arc or something?

OpenStudy (anonymous):

u want to express area in term of circumference?

OpenStudy (anonymous):

C =2piR so C/2Pi =r substituate and pi(C/2pi)^2

OpenStudy (anonymous):

yea, they just want you to make up a formula that you could use that has both of these equations in it

OpenStudy (anonymous):

C=2*pi*r -> pi*r= C/2 then plug in it into the equation for area